OpenStudy (anonymous):

2sin((2/3)x -(pi/6))

ganeshie8 (ganeshie8):

y = Asin(bx+c) amplitude = |A| period = 2pi/b phase shift = c

OpenStudy (anonymous):

A:2 Period: 2pi/(2/3) phaseshift: pi/6

ganeshie8 (ganeshie8):

Yes ! phase shift is pi/6 units to right

OpenStudy (anonymous):

awesome

ganeshie8 (ganeshie8):

looks u nailed these :)

OpenStudy (anonymous):

one last question for the night...

ganeshie8 (ganeshie8):

yes u may ask here

OpenStudy (anonymous):

i hope so i have a huge test tomorrow

OpenStudy (anonymous):

okies

ganeshie8 (ganeshie8):

u wil do great im sure

OpenStudy (anonymous):

y= cos ((pi/2)-x)

OpenStudy (anonymous):

i hope so !

ganeshie8 (ganeshie8):

we use this first : cos(-x) = cos(x)

ganeshie8 (ganeshie8):

y= cos ((pi/2)-x) is same as y= cos (x - (pi/2))

ganeshie8 (ganeshie8):

next do it exactly SAME as sin.

OpenStudy (anonymous):

A:1 pasheshift = -1 period :4

ganeshie8 (ganeshie8):

NO. you need to look at the thing attached to variable \(x\)

ganeshie8 (ganeshie8):

y= cos ((pi/2)-x) is same as y= cos (x - (pi/2)) ^^

ganeshie8 (ganeshie8):

A : 1 phaseshift : pi/2 rightside period : 2pi/1 = 2pi

ganeshie8 (ganeshie8):